a triangular garden has an area of 154 square feet. its height is 8 feet more than its base. find the measure of the base.
Answer by xcentaur(357)

area=154 sp feet=1/2bh
h=b+8
then area becomes:
154=b(b+8)/2
154*2=b(b+8)
308=b^2+8b
b^2+8b-308=0
(b-14)(b+22)=0
b=14,-22


length of base cannot be negative
hence b=14 feet


Hope this helps,
Prabhat